Grade 10Algebra

Hello Sir,

How many nos. are there between 100-1000 such that atleast one of the digits is 6??

Please answer.

Akshay Kandpal
14 Years agoGrade 10

3 Answers

Chetan Mandayam Nayakar
14 Years ago

Answer=9*10 +1=91

Krishna Pradyumna Mokshagundam
14 Years ago

no.s having 6 in their units place..... Ex:106,116,126,..... 206,216,226,....306,316...    = 10*9=90

no.s having 6 in tens place.... ex:160,260,360.....960  .....  =9

no.s having 6 in hundreds place.... 600 is the only number.  =1

Total numbers having atleast one six in their digits is 90+9+1=100

Nikhil panduranga
14 Years ago

between 100-1000 there are 900 "three" digit numbers  

there are a total of 10 nos 0-9 CASE 1 when no digit is 6

(8*8*7) Case 2 :all the possibilities i.e (9*8*7) Subtract them both u will have the answer............AM I RIGHT????